OpenStudy (anonymous):

Jules kicks a soccer ball off the ground and into the air with an initial velocity of 25 feet per second. Assume the starting height of the ball is 0 feet. Approximately what maximum height does the soccer ball reach?

OpenStudy (anonymous):

using the SUVAT equation v^2 = u^2 - 2gs, s = (u^2-v^2)/2g at max ht, v = 0, so s = 25^2/(2*32) ≈ 9.8 ft <-------

OpenStudy (anonymous):

dose this help?

OpenStudy (anonymous):

Yes. Thank you. @helper99